Use the definite integral to find the area between the​ x-axis and​ f(x) over the indicated interval. Check first to see if the graph crosses the​ x-axis in the given interval.
​f(x)=36-x^2​;​[​-1,​12]
The area between the​ x-axis and​ f(x) is
Please answer the area between the​ x-axis and​ f(x)

Answer by Alan3354(69443)

Use the definite integral to find the area between the​ x-axis and​ f(x) over the indicated interval. Check first to see if the graph crosses the​ x-axis in the given interval.
----
​f(x)=36-x^2​;​[​-1,​12]
\=======================
x^2 = 36 ---> x-intercepts at -6 and +6
-------------------
Int(36 - x^2) = g(x) = 36x - x^3/3 + C
Area from -1 to +6 = g(6) - g(-1)
Area = 216 - (-36 + 1/3) = 179 2/3 sq units
=================
Area from 6 to 12:
Area = g(12) - g(6) = 288 sq units ------> you can do the calculations.
-----
The area from 6 to 12 is below the x-axis, sometimes viewed as negative.
Subtracting would give a negative area, so add them.